OpenStudy (anonymous):

A compound that is composed of molybdenum (Mo) and oxygen (O) was produced in a lab by heating molybdenum over a Bunsen burner. The following data was collected: Mass of crucible: 38.26 g Mass of crucible and molybdenum: 39.52 g Mass of crucible and molybdenum oxide: 39.84 g Solve for the empirical formula of the compound, showing (or explaining in complete sentences) your calculations.

OpenStudy (anonymous):

Mass of molybdenum = 39.52-38.26 =1.26 g. Mass of molybdenum oxide = 39.84-38.26= 1.58 g Mass of oxygen = 1.58 - 1.26 = 0.32. No. of moles of molybdenum = 1.26 /96 = 0.0131 moles No. of moles of oxygen = 0.32/16 = 0.02 moles. molybdenum and oxygen are roughly in 1:2 molar ratio.So the compound is MoO2

OpenStudy (anonymous):

oh!! thanks!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

i got this wrong